OpenStudy (anonymous):

Find the equation of the line which cuts the x-axis at 5 and the y-axis at -2

OpenStudy (anonymous):

2 points are (5, 0) and (0, -2). The first is (x1, y1) and the second is (x2, y2), both in the form (x, y). First, get the slope from:\[\frac{ y _{1}- y _{2} }{ x _{1}-x _{2} } = m\]where "m" is the slope. Show your work, get the slope, and then we can go to the next step.

OpenStudy (anonymous):

How are you coming along with this?

OpenStudy (anonymous):

\[\frac{ 0-2 }{5-0 }=\frac{ -2 }{ 5 }\]

OpenStudy (anonymous):

oh it´s supposed to be 2 instead of -2

OpenStudy (anonymous):

Good! Now I'll get you the next step..

OpenStudy (anonymous):

Now that you have m = 2, the point-slope equation for the line is:\[y - y _{1} = m(x - x _{1})\]You can either leave the equation is this form (after making the substitutions using either point ) of rearrange into the standard or slope-intercept form of the line.

OpenStudy (anonymous):

\[\frac{ y-0 }{x-5 }=\frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

5(y-0)=2(x-5)

OpenStudy (anonymous):

5y-0=2x-10

OpenStudy (anonymous):

2x-5y=10

OpenStudy (anonymous):

Great work! Well done!

OpenStudy (anonymous):

Thanks for your help!!

OpenStudy (anonymous):

yw!

OpenStudy (anonymous):

Good work and good luck to you! If you like my work, hit "best response"